. A cylindrical rod has a length of 30 cm and a diameter of 2.4 cm. One end is kept in a fire at 500◦C while the other is held in cool water kept at 14◦C. What is the rate of heat conduction across the rod?
Answer #1

Conductive heat transfer can be expressed with Fourier's Law:

q = k A dT / s

where

q = heat transfer

A = heat transfer area

k = thermal conductivity of the material

dT = temperature gradient - difference - in the material

s = material thickness

Here A= 3.14*0.012*0.012=4.52*10^-4 m^2

d=0.3m

dT=486

Considering the rod is of iron, k=80

So q=80*4.52*10^-4*486/0.3

So q=58.57 W